# Large Language Models for Information Retrieval: A Survey

### v2

@@ 3.2.1 Inherent Knowledge of LLMs
LLMs are capable of storing knowledge within their parameters, making it a natural choice to capitalize on this knowledge for the purpose of query rewriting. As a pioneering work in LLM-based query rewriting, HyDE [92] generates a hypothetical document by LLMs according to the given query and then uses a dense retriever to retrieve relevant documents from the corpus that are relevant to the generated document. Query2doc [86] generates pseudo documents via prompting LLMs with few-shot demonstrations, and then expands the query with the generated pseudo document. Furthermore, the influence of different prompting methods and various model sizes on query rewriting is also investigated [93]. To better accommodate the frozen retriever and the LLM-based reader, a small language model is employed as the rewriter that is trained using reinforcement learning techniques with the rewards provided by the LLM-based reader [94]. It is worth noting that all of these works rely on the knowledge stored in LLMs rather than additional corpora. Despite this, experimental results demonstrate significant improvements compared to traditional query rewrite methods such as RM3 [87].
3.2.2 Inherent Knowledge of LLMs and Document Corpus
@@
Although LLMs exhibit remarkable capabilities, their lack of familiarity with specific domains can lead to the generation of hallucinatory or irrelevant queries. To address this issue, recent studies [95, 96, 97, 98] have proposed a hybrid approach that enhancing LLM-base query rewriting methods with document corpus.
Why incorporating a document corpus? The integration of a document corpus offers several notable advantages. First, it provides domain-specific knowledge by fine-tuning the query generation process for specific subject areas, enabling a targeted and specialized approach to IR. For instance, LLMs may not possess immediate comprehension of the query “ADHD”, but the examination of retrieved documents can offer valuable insights indicating its association with a medical condition referred to as “Attention-Deficit/Hyperactivity Disorder”. Second, it ensures the grounding of queries in reliable and verifiable knowledge by incorporating factual information extracted from the corpus. Third, it incorporates up-to-date concepts by enriching queries with contemporary information, surpassing the knowledge contained within LLMs, and ensuring enhanced information richness and timeliness. Finally, it supports relevance by utilizing relevant documents as a supplementary resource to refine the query generation process, effectively reducing the generation of irrelevant information and improving the production of contextually relevant outputs.
@@
How to incorporate a document corpus? In light of these advantages, various paradigms have been proposed to incorporate a document corpus into LLM-based query rewriting, which we summarize as follows.
@@
$\bullet$ Late fusion of LLM-only re-writing and traditional PRF retrieval results. Traditional PRF methods leverage relevant documents retrieved from a document corpus to rewrite queries, which restricts the query to the information contained in the target corpus. On the contrary, LLM-only re-writing methods provides external context not present in the corpus, which is more diverse. Both approaches have potential to independently enhance retrieval performance. Therefore, a straightforward strategy for combining them is using a weighted fusion method on retrieval results, which is demonstrated effective [98].
$\bullet$ Combining retrieved relevant documents in the prompts of LLMs. In the era of LLMs, incorporating instructions within the prompts is a flexible method for achieving specific functionalities. There are several studies that leverage this procedure. For the query intent classification task, QUILL [96] demonstrates that retrieval augmentation of queries provides LLMs with meaningful supportive context, which leads to better understanding. It proposes a two-stage distill pipeline to transfer the knowledge in LLMs to small models in order to rewrite queries more efficiently. LameR [99] proposes a retrieve-rewrite-retrieve framework. It takes in BM25 as the retriever without using any annotated query-document pairs, which shows the great potential of LLMs in zero-shot retrieval. Besides, there are no strict rules about whether rewriters should appear before retrievers. InteR [97] proposes a framework supporting multi-turn interaction between search engines and LLMs. This setup enables search engines to expand queries with LLM-generated knowledge, while simultaneously permitting LLMs to refine prompt formulations with relevant documents provided by search engines.
@@
$\bullet$ Enhancing factuality of generative relevance feedback (GRF) by pseudo relevance feedback (PRF). Although generative documents are often relevant and diverse, they exhibit hallucinatory characteristics. In contrast, traditional documents are generally regarded as reliable sources of factual information. Motivated by this observation, GRM [95] proposes a novel technique known as relevance-aware sample estimation (RASE). RASE leverages relevant documents retrieved from the collection to assign weights to generated documents. In this way, GRM ensures that relevance feedback is not only diverse but also maintain a high degree of factuality.
3.3 Generation Methodology
There are three primary paradigms employed to leverage LLMs for the task of query rewriting. The most prevalent paradigm is known as prompting methods, in which a specific prompt or instruction is designed to guide the language model in generating the desired output. Prompting methods offer researchers flexibility and interpretability over the rewriting process. In addition to prompting methods, fine-tuning LLMs in domain-specific areas and knowledge distillation are also effective methods. LLM contains world knowledge, which is general but may be ineffective for specific domains. Fine-tuning entails training a pre-trained LLM on a specific dataset or task in order to enhance its domain-specific effectiveness. Retrieval augmentation causes increased complexity of LLM inference. Knowledge distillation alleviates this problem by transferring the knowledge of LLMs to some lightweight models. In the following section, we will introduce these three methods in detail.
3.3.1 Prompting Methods
@@
TABLE I: Examples of different prompting methods in query rewriting.
@@
Chine-of-Thought
CoT [93] Answer the following query: {#Query} Give the rationale before answering
CoT/PRF [93] Answer the following query based on the context:
@@
Prompting in LLMs refers to the technique of providing a specific instruction or context to guide the model’s generation of text. The form of prompts can vary, ranging from questions and incomplete sentences to specific instructions. The prompt serves as a conditioning signal and influences the language generation process of the model. Consequently, the effectiveness of prompting is dependent on the quality and design of the prompt.
The prevailing prompting strategies generally encompass three categories: zero-shot prompting, few-shot prompting, and chain-of-thought (CoT) prompting [45].
@@
$\bullet$ Zero-shot prompting. Zero-shot prompting involves instructing the model to generate text on a specific topic without any prior exposure to training examples in that domain or topic. The model relies on its pre-existing knowledge and language understanding to generate coherent and contextually relevant expanded terms for original queries. Experiments show that zero-shot prompting is a simple but effective method for query rewriting [100, 101, 98, 99, 97, 93].
@@
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [100, 101, 93, 86]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to the specific task or domain at hand. Query2Doc [86] prompts LLMs to write a document that answers the query with some demo query-document pairs provided by the ranking dataset, such as MSMARCO [102] and NQ [103]. This work experiment on a single prompt. To further study the impact of different prompt designing,  [93] explores eight different prompts, such as promting LLMs to generate query expansion terms instead of entire pseudo documents, CoT prompting and so on. There are some illustrative prompts in Table I. This work conducts more experiments than Query2Doc, but experiments show that this prompt is not as effective as that in Query2Doc.
@@
$\bullet$ Chain-of-thought prompting. CoT prompting [45] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [100, 93]. In conversational search, the process of query re-writing is multi-turn, which means query should be refined step-by-step with the interaction between search engine and users. This process is naturally coincided with CoT process. As shown in 4, engineers conduct the CoT process through adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one-round in query re-writing. So CoT could only be accomplished in a simple and coarse way. For example, as shown in Table I, researchers add “Give the rationale before answering” in the instructions to prompt LLMs think deeply [93].
3.3.2 Fine-tuning Methods
Fine-tuning is another effective and prevalent paradigm to let LLMs adapt to the specific domain better. The process of fine-tuning typically involves taking a pre-trained LM, such as GPT-3, and further training it on a dataset that is specific to the target domain. The dataset may consist of queries, rewrites, and associated labels, or it may be generated through a combination of human expertise and data augmentation techniques. During fine-tuning, the model’s parameters are adjusted to optimize its performance on the domain-specific task.
By training the LLM on domain-specific data, it can learn domain-specific patterns, terminology, and context, which can enhance its ability to generate high-quality query rewrites. However, fine-tuning LLMs could be quite expansive. There are several studies trying to fine-tune LLM on query rewriting tasks. A typical study fine-tunes LLMs to generate relevant documents of the given query [101], and combines query with the generative documents as new query. Concretely, it fine-tunes FiD [104] (a variant of T5) with 770M and 3B parameters on downstream datasets such as NQ [103] and FEVER [105]. Experimental results show that fine-tuning the 770M FiD model with an input length of 550 tokens can be easily performed on a single Tesla V100 32GB GPU, while training the 3B T5 model requires a larger cluster consisting of eight Tesla V100 or A100 GPUs.
3.3.3 Knowledge Distillation Methods
@@
Although LLM-based methods have demonstrated significant improvements in query rewriting tasks, their practical implementation for online deployment is hindered by the substantial latency caused by the computational requirements of LLMs. To mitigate this challenge, knowledge distillation has emerged as a prominent technique in the industry. In the QUILL [96] framework, a two-stage distillation method is proposed. This approach entails utilizing a retrieval-augmented LLM as the professor model, a vanilla LLM as the teacher model, and a lightweight BERT model as the student model. The professor model is trained on two extensive datasets, namely Orcas-I [106] and EComm [96], which are specifically curated for query intent understanding. Subsequently, a two-stage distillation process is employed to transfer knowledge from the professor model to the teacher model, followed by knowledge transfer from the teacher model to the student model. Empirical findings demonstrate that this knowledge distillation methodology surpasses the simple scaling up of model size from base to XXL, resulting in even more substantial improvements. In a recently proposed “rewrite-retrieve-read” framework [94], an LLM is first used to rewrite the queries by prompting, followed by a retrieval-augmented reading process. To improve framework effectiveness, a trainable rewriter, implemented as a small language model, is incorporated to further adapt search queries to align with both the frozen retriever and the LLM reader’s requirements. The rewriter’s refinement involves a two-step training process. Initially, supervised warm-up training is conducted using pseudo data. Subsequently, the retrieve-then-read pipeline is described as a reinforcement learning scenario, with the rewriter’s training acting as a policy model to maximize pipeline performance rewards.
@@
HyDE [96] Ad-hoc LLMs Prompting
Jagerman et al. [93] Ad-hoc LLMs Prompting
Query2Doc [86] Ad-hoc LLMs Prompting
Ma et al. [94] Ad-hoc LLMs Finetuning
@@ 3.4 Limitations
Experiments across various query rewrite methods consistently highlight the positive impact of increasing the number of generated documents or answers on retrieval performance [99]. While candidate documents contribute to presenting supplementary contexts of the original query, most existing works emphasize the original query’s prominence through repetition or alternative strategies [99, 97, 93]. According to our knowledge, there is a lack of dedicated metrics designed for evaluating query rewriting. Existing models heavily rely on downstream retrieval tasks to assess the effectiveness of query rewriters. How to directly determine whether the rewritten queries reflect human intent and effectively serve specific tasks remains unresolved.
@@ 4.1.2 Training Data Augmentation
How to apply LLMs for data augmentation? In the scenario of information retrieval, it is easy to collect numerous documents. However, the challenging and costly task lies in gathering real user queries and labeling the relevant documents accordingly. Considering the strong text generation capability of LLMs, many researchers [107, 108] suggest using LLM-driven processes to create pseudo queries or relevance labels based on existing collections. These approaches facilitate the construction of relevant query-document pairs, enlarging the training data for retrieval models. According to the type of generated data, there are two mainstream approaches that complement the LLM-based data augmentation for retrieval models, i.e., pseudo query generation and relevance label generation. Their frameworks are visualized in Figure 5. Next, we will give an overview of the related studies.
$\bullet$ Pseudo query generation. Given the abundance of documents, a straightforward idea is to use LLMs for generating their corresponding pseudo queries. One such illustration is presented by inPairs [107], which leverages the in-context learning capability of GPT-3. This method employs a collection of query-document pairs as demonstrations. These pairs are combined with a document and presented as input to GPT-3, which subsequently generates possible relevant queries for the given document. By combining the same demonstration with various documents, it is easy to create a vast pool of synthetic training samples and support the fine-tuning of retrievers on specific target domains. To enhance the reliability of these synthetic samples, a fine-tuned model (e.g., a monoT5-3B model fine-tuned on MSMARCO [108]) is employed to filter the generated queries. Only the top pairs with the highest estimated relevance scores are kept for training. This “generating-then-filtering” paradigm can be conducted iteratively in a round-trip filtering manner, i.e., by first fine-tuning a retriever on the generated samples and then filtering the generated samples using this retriever. Repeating these EM-like steps until convergence can produce high-quality training sets [109]. Furthermore, by adjusting the prompt given to LLMs, they can generate queries of different types. This capability allows for a more accurate simulation of real queries with various patterns [110].

### v3

@@ 3.2.1 LLM-only methods
LLMs are capable of storing knowledge within their parameters, making it a natural choice to capitalize on this knowledge for the purpose of query rewriting. As a pioneering work in LLM-based query rewriting, HyDE [101] generates a hypothetical document by LLMs according to the given query and then uses a dense retriever to retrieve relevant documents from the corpus that are relevant to the generated document. Query2doc [86] generates pseudo documents via prompting LLMs with few-shot demonstrations, and then expands the query with the generated pseudo document. Furthermore, the influence of different prompting methods and various model sizes on query rewriting has also been investigated [102]. To better accommodate the frozen retriever and the LLM-based reader, a small language model is employed as the rewriter that is trained using reinforcement learning techniques with the rewards provided by the LLM-based reader [100]. GFF [106] presents a “Generate, Filter, and Fuse” method for query expansion. It employs an LLM to create a set of related keywords via a reasoning chain. Then, a self-consistency filter is used to identify the most important keywords, which are concatenated with the original queries for the downstream reranking task.
It is worth noting that though the designs of these methods are different, all of them rely on the world knowledge stored in LLMs without additional corpora.
3.2.2 Corpus-enhanced LLM-based methods
@@
Although LLMs exhibit remarkable capabilities, the lack of domain-specific knowledge may lead to the generation of hallucinatory or irrelevant queries. To address this issue, recent studies [96, 97, 98, 99] have proposed a hybrid approach that enhances LLM-based query rewriting methods with an external document corpus.
Why incorporate a document corpus? The integration of a document corpus offers several notable advantages. Firstly, it boosts relevance by using relevant documents to refine query generation, reducing irrelevant content and improving contextually appropriate outputs. Second, enhancing LLMs with up-to-date information and specialized knowledge in specific fields enables them to effectively deal with queries that are both current and specific to certain domains.
@@
How to incorporate a document corpus? Thanks to the flexibility of LLMs, various paradigms have been proposed to incorporate a document corpus into LLM-based query rewriting, which can be summarized as follows.
@@
$\bullet$ Late fusion of LLM-based re-writing and pseudo relevance feedback (PRF) retrieval results. Traditional PRF methods leverage relevant documents retrieved from a document corpus to rewrite queries, which restricts the query to the information contained in the target corpus. On the contrary, LLM-based rewriting methods provide external context not present in the corpus, which is more diverse. Both approaches have the potential to independently enhance retrieval performance. Therefore, a straightforward strategy for combining them is using a weighted fusion method for retrieval results [99].
$\bullet$ Combining retrieved relevant documents in the prompts of LLMs. In the era of LLMs, incorporating instructions within the prompts is the most flexible method for achieving specific functionalities. QUILL [97] and CAR [107] illustrate how retrieval augmentation of queries can provide LLMs with context that significantly enhances query understanding. LameR [108] takes this further by using LLM expansion to improve the simple BM25 retriever, introducing a retrieve-rewrite-retrieve framework. Experimental results reveal that even basic term-based retrievers can achieve comparable performance when paired with LLM-based rewriters. Additionally, InteR [98] proposes a multi-turn interaction framework between search engines and LLMs. This enables search engines to expand queries using LLM-generated insights, while LLMs refine prompts using relevant documents sourced from the search engines.
@@
$\bullet$ Enhancing factuality of generative relevance feedback (GRF) by pseudo relevance feedback (PRF). Although generative documents are often relevant and diverse, they exhibit hallucinatory characteristics. In contrast, traditional documents are generally regarded as reliable sources of factual information. Motivated by this observation, GRM [96] proposes a novel technique known as relevance-aware sample estimation (RASE). RASE leverages relevant documents retrieved from the collection to assign weights to generated documents. In this way, GRM ensures that relevance feedback is not only diverse but also maintains a high degree of factuality.
3.3 Rewriting Approaches
There are three main approaches used for leveraging LLMs in query rewriting: prompting methods, fine-tuning, and knowledge distillation. Prompting methods involve using specific prompts to direct LLM output, providing flexibility and interpretability. Fine-tuning adjusts pre-trained LLMs on specific datasets or tasks to improve domain-specific performance, mitigating the general nature of LLM world knowledge. Knowledge distillation, on the other hand, transfers LLM knowledge to lightweight models, simplifying the complexity associated with retrieval augmentation. In the following section, we will introduce these three methods in detail.
3.3.1 Prompting
@@
TABLE I: Partial Examples of different prompting methods in query rewriting.
@@
Chain-of-Thought
CoT [102] Answer the following query based on the context:
@@
Prompting in LLMs refers to the technique of providing a specific instruction or context to guide the model’s generation of text. The prompt serves as a conditioning signal and influences the language generation process of the model. Existing prompting strategies can be roughly categorized into three groups: zero-shot prompting, few-shot prompting, and chain-of-thought (CoT) prompting [45].
@@
$\bullet$ Zero-shot prompting. Zero-shot prompting involves instructing the model to generate texts on a specific topic without any prior exposure to training examples in that domain or topic. The model relies on its pre-existing knowledge and language understanding to generate coherent and contextually relevant expanded terms for original queries. Experiments show that zero-shot prompting is a simple yet effective method for query rewriting [109, 110, 99, 108, 98, 102].
@@
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [109, 110, 102, 86]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to the specific task or domain at hand. Query2Doc [86] prompts LLMs to write a document that answers the query with some demo query-document pairs provided by the ranking dataset, such as MSMARCO [111] and NQ [112]. This work experiments with a single prompt. To further study the impact of different prompt designing, recent works [102] have explored eight different prompts, such as prompting LLMs to generate query expansion terms instead of entire pseudo documents and CoT prompting. There are some illustrative prompts in Table I. This work conducts more experiments than Query2Doc, but the results show that the proposed prompt is less effective than Query2Doc.
@@
$\bullet$ Chain-of-thought prompting. CoT prompting [45] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [109, 102]. In conversational search, the process of query re-writing is multi-turn, which means queries should be refined step-by-step with the interaction between search engines and users. This process is naturally coincided with CoT process. As shown in 4, users can conduct the CoT process through adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one-round in query re-writing, so CoT could only be accomplished in a simple and coarse way. For example, as shown in Table I, researchers add “Give the rationale before answering” in the instructions to prompt LLMs think deeply [102].
3.3.2 Fine-tuning
Fine-tuning is an effective approach for adapting LLMs to specific domains. This process usually starts with a pre-trained language model, like GPT-3, which is then further trained on a dataset tailored to the target domain. This domain-specific training enables the LLM to learn unique patterns, terminology, and context relevant to the domain, which is able to improve its capacity to produce high-quality query rewrites.
BEQUE [113] leverages LLMs for rewriting queries in e-commerce product searches. It designs three Supervised Fine-Tuning (SFT) tasks: quality classification of e-commerce query rewrites, product title prediction, and CoT query rewriting. To our knowledge, it is the first model to directly fine-tune LLMs, including ChatGLM [114, 68], ChatGLM2.0 [114, 68], Baichuan [115], and Qwen [116], specifically for the query rewriting task. After the SFT stage, BEQUE uses an offline system to gather feedback on the rewrites and further aligns the rewriters with e-commerce search objectives through an object alignment stage. Online A/B testing demonstrates the effectiveness of the method.
3.3.3 Knowledge Distillation
@@
Although LLM-based methods have demonstrated significant improvements in query rewriting tasks, their practical implementation for online deployment is hindered by the substantial latency caused by the computational requirements of LLMs. To address this challenge, knowledge distillation has emerged as a prominent technique in the industry. In the QUILL [97] framework, a two-stage distillation method is proposed. This approach entails utilizing a retrieval-augmented LLM as the professor model, a vanilla LLM as the teacher model, and a lightweight BERT model as the student model. The professor model is trained on two extensive datasets, namely Orcas-I [117] and EComm [97], which are specifically curated for query intent understanding. Subsequently, a two-stage distillation process is employed to transfer knowledge from the professor model to the teacher model, followed by knowledge transfer from the teacher model to the student model. Empirical findings demonstrate that this knowledge distillation methodology surpasses the simple scaling up of model size from base to XXL, resulting in even more substantial improvements. In a recently proposed “rewrite-retrieve-read” framework [100], an LLM is first used to rewrite the queries by prompting, followed by a retrieval-augmented reading process. To improve framework effectiveness, a trainable rewriter, implemented as a small language model, is incorporated to further adapt search queries to align with both the frozen retriever and the LLM reader’s requirements. The rewriter’s refinement involves a two-step training process. Initially, supervised warm-up training is conducted using pseudo data. Then, the retrieve-then-read pipeline is described as a reinforcement learning scenario, with the rewriter’s training acting as a policy model to maximize pipeline performance rewards.
@@
HyDE [97] Ad-hoc LLMs Prompting
Jagerman et al. [102] Ad-hoc LLMs Prompting
Query2Doc [86] Ad-hoc LLMs Prompting
Ma et al. [100] Ad-hoc LLMs Finetuning
PromptCase [103] Ad-hoc LLMs Prompting
@@
CAR [107] Ad-hoc LLMs + Docs Prompting
@@
CONVERSER [105] Conversational LLMs Prompting
Ye et al. [104] Conversational LLMs Prompting
@@ 3.4 Limitations
While LLMs offer promising capabilities for query rewriting, they also meet several challenges. Here, we outline two main limitations of LLM-based query rewriters.
3.4.1 Concept Drifts
When using LLMs for query rewriting, they may introduce unrelated information, known as concept drift, due to their extensive knowledge base and tendency to produce detailed and redundant content. While this can enrich the query, it also risks generating irrelevant or off-target results.
This phenomenon has been reported in several studies [118, 107, 113] These studies highlight the need for a balanced approach in LLM-based query rewriting, ensuring that the essence and focus of the original query are maintained while leveraging the LLM’s ability to enhance and clarify the query. This balance is crucial for effective search and IR applications.
3.4.2 Correlation between Retrieval Performance and Expansion Effects
Recently, a comprehensive study [119] conduct experiments on various expansion techniques and downstream ranking models, which reveals a notable negative correlation between retriever performance and the benefits of expansion. Specifically, while expansion tends to enhance the scores of weaker models, it generally hurts stronger models. This observation suggests a strategic approach: employ expansions with weaker models or in scenarios where the target dataset substantially differs in format from the training corpus. In other cases, it is advisable to avoid expansions to maintain clarity of the relevance signal.
@@ 4.1.1 Search Data Refinement
Ma et al. [121] 0-2 Alpaca-LLaMA & tk-Instruct Relevant query - Fixed
@@ 4.1.2 Training Data Augmentation
How to apply LLMs for data augmentation? In the scenario of IR, it is easy to collect numerous documents. However, the challenging and costly task lies in gathering real user queries and labeling the relevant documents accordingly. Considering the strong text generation capability of LLMs, many researchers [120, 122] suggest using LLM-driven processes to create pseudo queries or relevance labels based on existing collections. These approaches facilitate the construction of relevant query-document pairs, enlarging the training data for retrieval models. According to the type of generated data, there are two mainstream approaches that complement the LLM-based data augmentation for retrieval models, i.e., pseudo query generation and relevance label generation. Their frameworks are visualized in Figure 5. Next, we will give an overview of the related studies.
$\bullet$ Pseudo query generation. Given the abundance of documents, a straightforward idea is to use LLMs for generating their corresponding pseudo queries. One such illustration is presented by inPairs [120], which leverages the in-context learning capability of GPT-3. This method employs a collection of query-document pairs as demonstrations. These pairs are combined with a document and presented as input to GPT-3, which subsequently generates possible relevant queries for the given document. By combining the same demonstration with various documents, it is easy to create a vast pool of synthetic training samples and support the fine-tuning of retrievers on specific target domains. Recent studies [121] have also leveraged open-sourced LLMs, such as Alpaca-LLaMA and tk-Instruct, to produce sufficient pseudo queries and applied curriculum learning to pre-train dense retrievers. To enhance the reliability of these synthetic samples, a fine-tuned model (e.g., a monoT5-3B model fine-tuned on MSMARCO [122]) is employed to filter the generated queries. Only the top pairs with the highest estimated relevance scores are kept for training. This “generating-then-filtering” paradigm can be conducted iteratively in a round-trip filtering manner, i.e., by first fine-tuning a retriever on the generated samples and then filtering the generated samples using this retriever. Repeating these EM-like steps until convergence can produce high-quality training sets [123]. Furthermore, by adjusting the prompt given to LLMs, they can generate queries of different types. This capability allows for a more accurate simulation of real queries with various patterns [124].
@@
In addition to the above studies, pseudo query generation methods are also introduced in other application scenarios, such as conversational dense retrieval [105] and multilingual dense retrieval [129].